### v1

@@ Abstract
The safety alignment of large language models (LLMs) can be circumvented through adversarially crafted inputs, yet the mechanisms by which these attacks bypass safety barriers remain poorly understood. Prior work suggests that a single refusal direction in the model’s activation space determines whether an LLM refuses a request. In this study, we propose a novel gradient-based approach to representation engineering and use it to identify refusal directions. Contrary to prior work, we uncover multiple independent directions and even multi-dimensional concept cones that mediate refusal. Moreover, we show that orthogonality alone does not imply independence under intervention, motivating the notion of representational independence that accounts for both linear and non-linear effects. Using this framework, we identify mechanistically independent refusal directions. We show that refusal mechanisms in LLMs are governed by complex spatial structures and identify functionally independent directions, confirming that multiple distinct mechanisms drive refusal behavior. Our gradient-based approach uncovers these mechanisms and can further serve as a foundation for future work on understanding LLMs.1
@@ 1 Introduction
The breakthrough of scaling large language models (LLMs) has led to an unprecedented leap in capabilities, driving widespread real-world adoption OpenAI (2022). However, these advancements also introduce serious risks. As artificial intelligence becomes more powerful, it can be misused for harmful purposes, such as attacking critical
@@
infrastructure or spreading misinformation. Ensuring that these models remain aligned with human values has become a crucial research challenge Liu et al. (2023); Schwinn et al. (2025). Despite significant progress, LLMs, like all machine learning models, remain vulnerable to adversarial attacks that can bypass alignment mechanisms and induce harmful outputs Szegedy et al. (2014); Carlini et al. (2024).
Recent work in interpretability has provided valuable insights into how LLMs encode and process information Nanda et al. (2024); Wang et al. (2022); Cunningham et al. (2023); Heinzerling & Inui (2024). Prior studies Belrose et al. (2023); Gurnee & Tegmark (2023); Marks & Tegmark (2024) suggest that concepts—ranging from simple to complex—are often encoded linearly in the model’s residual stream. Methods such as representation engineering Zou et al. (2023a) allow researchers to use input prompts to analyze model behavior by extracting and manipulating such concepts. However, the mechanisms enabling adversarial jailbreaks that bypass alignment safeguards remain poorly understood. Some evidence suggests that refusals to harmful queries are mediated by a single “refusal direction” in activation space Arditi et al. (2024), and that jailbreaks rely on manipulating this direction Yu et al. (2024), yet these assumptions require further examination.
In this work, we go beyond extracting concepts using common input prompt methods by introducing a novel gradient-based approach to representation engineering which we use to investigate the mechanisms underlying refusal behavior in LLMs. We extract refusal-mediating directions more effectively, improving both precision and control while minimizing unintended side effects, which we demonstrate in Section 4. Unlike prior work that assumes model refusal is controlled by a single linear direction, we show in Section 5 that there exist multi-dimensional polyhedral cones which contain infinite refusal directions; we show an illustrative example in Figure 1. To further characterize refusal mechanisms in language models, we introduce representational independence, a criterion for identifying directions that remain mutually unaffected under intervention, capturing both linear and non-linear dependencies across layers. In Section 6, we demonstrate that even under this strict notion of independence, multiple complementary refusal directions exist.
@@ 2 Background
Notation.
Let $f:{\mathbb{T}}^{N_{\textnormal{seq}}}\to\Delta^{N_{\textnormal{seq}}\times|{\mathbb{T}}|}$ denote a language model, where $\Delta^{|{\mathbb{T}}|}$ is the probability simplex over vocabulary ${\mathbb{T}}$. Given a prompt ${p}=(t_{1},\dots,t_{N_{\textnormal{seq}}})\in{\mathbb{T}}^{N_{\textnormal{seq}}}$ consisting of tokens $t_{i}$, each token is first embedded: $x_{i}^{(0)}=\textsc{Embed}(t_{i})$. The model then processes the token sequence through $L$ layers, where at each layer $l=1,\dots,L$ and token position $i$ the following transformation is applied:
@@
Extracting refusal directions. Paired prompts of harmful and harmless requests allow the extraction of a directional feature from the model’s residual stream as shown by prior work Panickssery et al. (2024); Bolukbasi et al. (2016); Burns et al. (2024). Recent studies obtain this direction by computing the difference-in-means (DIM) Panickssery et al. (2024); Arditi et al. (2024); Stolfo et al. (2024) between model representations on datasets of harmful prompts $\mathcal{D}_{\text{harm}}$ and harmless prompts $\mathcal{D}_{\text{good}}$:
@@
Here, ${\bm{x}}_{i}^{(l)}({p})$ represents the residual stream activations at position $i$, layer $l$ for input prompt ${p}$.
Adversarial steering attacks. The extracted harmfulness direction can be used to manipulate the model’s refusal behavior. With white-box access, an attacker can prompt the model with harmful queries and suppress activations in the harmfulness direction, thereby reducing the model’s probability of refusal. This can be done through directional ablation of ${{\bm{r}}}$ (where ${\hat{{{\bm{r}}}}}$ denotes the unit vector) Zou et al. (2023a):
@@
which subtracts a scaled ${{\bm{r}}}$ from the residual stream.We follow common practice to apply both operations across all token positions and ablation across all layers while doing subtraction only at a single layer.
@@ 3 Related Work
Adversarial attacks for LLMs. Many studies have explored hand-crafted adversarial techniques, such as persona modulation Shah et al. (2023), language modifications Zhu et al. (2023), or prompt engineering using repetitions and persuasive phrasing Rao et al. (2024). Other works take a more systematic approach, employing techniques like genetic algorithms and random search Chen et al. (2024), discrete optimization over input tokens Zou et al. (2023b), or gradient-based methods to identify high-impact perturbations Geisler et al. (2024). While identifying these vulnerabilities enables adversarial fine-tuning Xhonneux et al. (2024) or improved training through Reinforcement Learning with Human Feedback (RLHF), recent works suggest that robustness remains a challenge Zou et al. (2023a); Schwinn et al. (2024); Geisler et al. (2024); Scholten et al. (2025).
@@
Interpretability of LLMs. A parallel line of research focuses on understanding the internal mechanisms of LLMs, as their natural language outputs provide a unique opportunity to link internal states to interpretable behaviors. Interpretability research has led to the identification of various “features”—concepts represented by distinct activation patterns Cunningham et al. (2023)—as well as “circuits”, which are subnetworks that implement a specific function or behavior. Prominent examples are backup circuits Nanda et al. (2024) and information mover circuits Wang et al. (2022). Many interpretability insights rely on extracting features using paired inputs with opposing semantics Burns et al. (2024) and then manipulating residual stream activations to elicit specific behaviors Panickssery et al. (2024). Representation engineering, as proposed by Zou et al. (2023a), investigates the linear representation of concepts such as truthfulness, honesty, and fairness in LLMs. The effectiveness of these methods supports the hypothesis that many features are encoded linearly in LLMs Marks & Tegmark (2024). These insights allow researchers to pinpoint and manipulate concept representations or specific circuits, enabling targeted debugging of behaviors, mitigating biases, and advancing safer, more reliable AI systems.
@@
Understanding Refusal Mechanisms. Recent research has focused on understanding the mechanisms underlying refusal behaviors in LLMs. For example, removing safety-critical neurons has been shown to decrease robustness Wei et al. (2024); Li et al. (2024b). Zheng et al. (2024) demonstrate that adding explicit safety prompts shifts the internal representation along a harmfulness direction. O’Brien et al. (2024) propose to use sparse autoencoders to identify latent features that mediate refusal. The most relevant work to ours is Arditi et al. (2024), which builds on Zou et al. (2023a) and examines the representation of refusal in LLMs. Their work suggests that a single direction a model’s activation space determines whether the model accepts or refuses a request. We challenge this notion by showing that refusal is mediated through more nuanced mechanisms.
4 Gradient-based Refusal Directions
@@
Research Question: Can gradient-based representation engineering identify refusal directions?
@@
To investigate the refusal mechanisms in language models, we propose a gradient-based algorithm that identifies directions controlling refusal in the model’s activation space. We refer to it as Refusal Direction Optimization (RDO). Unlike prior approaches that extract refusal directions using paired prompts of harmless and harmful instructions Arditi et al. (2024), our method leverages gradients to find better directions instead of solely relying on model activations. Similar to Park et al. (2023), we define two key properties for refusal directions:
@@
  • Monotonic Scaling: when using the direction for activation addition/subtraction
    $\smash{\check{{\bm{x}}}^{(l)}_{i}={\bm{x}}^{(l)}_{i}+\alpha\cdot{{\bm{r}}}}$, the model’s probability of refusing instructions should scale monotonically with $\alpha$.
@@
We can encode the desired refusal properties into loss functions, allowing us to find corresponding refusal vectors ${{\bm{r}}}$ using gradient descent. For the monotonic scaling property, we train the model to refuse harmless instructions ${{p}_{\text{safe}}}$ when running the model $f$ with a modified forward pass $f_{\text{add}({{\bm{r}}},l)}$ in which we add ${{\bm{r}}}$ to the activations at layer $l$. We minimize the cross-entropy between the model output and target refusal response ${t_{\text{refusal}}}$. For the surgical ablation property, we similarly compute the cross-entropy between a harmful response target ${t_{\text{answer}}}$ and the output of a modified forward pass $f_{\text{ablate}({{\bm{r}}})}$ to make the model respond to harmful instructions. A key strength of our gradient-based approach is the ability to control any predefined objective and thus we can control the extent to which other concepts are affected during interventions. For this, we use a retain loss based on the Kullback-Leibler (KL) divergence to ensure that directional ablation of ${{\bm{r}}}$ on harmless instructions does not change the model’s output over a target response ${t_{\text{retain}}}$. Algorithm 1 shows the full training procedure for our refusal directions.
@@
Setup. We construct a dataset of harmless and harmful prompts from the alpaca Taori et al. (2023) and salad-bench Li et al. (2024a) datasets (see Section A.1). An important consideration for our algorithm is the choice of targets ${t_{\text{answer}}}$ and ${t_{\text{refusal}}}$. Generally, language models differ in their refusal and response styles, which is why we use model-specific targets rather than generating them via uncensored LLMs as in Zou et al. (2024). Specifically, we use the DIM refusal direction to generate our targets, though any effective attack can work. For the harmful answers ${t_{\text{answer}}}$, we ablate the DIM direction and generate 30 tokens. Similarly, we use activation addition on harmless instructions to produce refusal targets ${t_{\text{refusal}}}$. For helpful answers on harmless instructions that should be retained ${t_{\text{retain}}}$, we generate 29 tokens without intervention. The retain loss $\mathcal{L}_{\text{retain}}$ is applied over the last 30 tokens, such that the last token of the model’s chat template is included. We detail hyperparameters and implementation in Appendix A.
@@
Evaluation. We evaluate our method by training a refusal direction on various models from the Gemma 2 Team et al. (2024), Qwen2.5 Yang et al. (2024), and Llama-3 Dubey et al. (2024) families and compare against the DIM direction for which we use the same setup as Arditi et al. (2024) but with our expanded dataset. For a fair comparison, we train the refusal direction at the same layer that the DIM direction is extracted from, and during activation addition/subtraction set the scaling coefficient $\alpha$ to the norm of the DIM direction. We evaluate the jailbreak Attack Success Rate (ASR) on JailbreakBench Chao et al. (2024) using the StrongREJECT fine-tuned judge Souly et al. (2024). For inducing refusal via activation addition, we test 128 harmless instructions sampled from alpaca using substring matching of common refusal phrases. Model completions for evaluation are generated using greedy decoding with a maximum generation length of 512 tokens.
@@
Does the direction mediate refusal? In Figure 2, we show that for jailbreaking, our approach is competitive when using directional ablation and, on average, outperforms DIM when subtracting the refusal direction. Notably, despite not being explicitly optimized for subtraction-based attacks, our direction naturally generalizes to this setting. Figure 9 shows that adding the refusal direction to harmless inputs induces refusal more effectively with RDO than with DIM, further indicating that our method manipulates refusal more effectively.
@@
Is the direction more precise? To measure the side effects when intervening with the directions we track benchmark performance. Arditi et al. (2024) show that directional ablation with the DIM direction tends to have little impact on benchmark performance, except for TruthfulQA Lin et al. (2021). In Table 1, we show that RDO impacts TruthfulQA performance much less severely, reducing the error by $40\%$ on average.
@@
Is our method versatile? Hyperparameter tuning of the retain loss weight $\lambda_{\text{ret}}$ in Algorithm 1 allows for balancing between attack success and side effects (see Appendix B). We observe that for many models—especially those in the Qwen 2.5 family—for the majority of estimated DIM directions, the side-effects are too high, rendering it an unsuccessful attack (see Figure 16). Our method is more flexible than previous work as we can choose the target layer freely while limiting side effects through the retain loss (if possible).
Key Takeaways. Our RDO yields more effective refusal directions with fewer side effects, establishing that gradient-based representation engineering is an effective approach for extracting meaningful directions, while allowing for more modeling freedom such as incorporating side constraints.

### v2

@@ Abstract
The safety alignment of large language models (LLMs) can be circumvented through adversarially crafted inputs, yet the mechanisms by which these attacks bypass safety barriers remain poorly understood. Prior work suggests that a single refusal direction in the model’s activation space determines whether an LLM refuses a request. In this study, we propose a novel gradient-based approach to representation engineering and use it to identify refusal directions. Contrary to prior work, we uncover multiple independent directions and even multi-dimensional concept cones that mediate refusal. Moreover, we show that orthogonality alone does not imply independence under intervention, motivating the notion of representational independence that accounts for both linear and non-linear effects. Using this framework, we identify mechanistically independent refusal directions. We show that refusal mechanisms in LLMs are governed by complex spatial structures and identify functionally independent directions, confirming that multiple distinct mechanisms drive refusal behavior. Our gradient-based approach uncovers these mechanisms and can further serve as a foundation for future work on understanding LLMs.
Machine Learning, ICML
@@ 1 Introduction
The breakthrough of scaling large language models (LLMs) has led to an unprecedented leap in capabilities, driving widespread real-world adoption (OpenAI, 2022). However,
@@
these advancements also introduce serious risks. As artificial intelligence becomes more powerful, it can be misused for harmful purposes, such as attacking critical infrastructure or spreading misinformation. Ensuring that these models remain aligned with human values has become a crucial research challenge (Liu et al., 2023; Schwinn et al., 2025). Despite significant progress, LLMs, like all machine learning models, remain vulnerable to adversarial attacks that can bypass alignment mechanisms and induce harmful outputs (Szegedy et al., 2014; Carlini et al., 2024).
Recent work in interpretability has provided valuable insights into how LLMs encode and process information (Nanda et al., 2024; Wang et al., 2022; Cunningham et al., 2023; Heinzerling & Inui, 2024). Prior studies (Belrose et al., 2023; Gurnee & Tegmark, 2023; Marks & Tegmark, 2024) suggest that concepts—ranging from simple to complex—are often encoded linearly in the model’s residual stream. Methods such as representation engineering (Zou et al., 2023a) allow researchers to use input prompts to analyze model behavior by extracting and manipulating such concepts. However, the mechanisms enabling adversarial jailbreaks that bypass alignment safeguards remain poorly understood. Some evidence suggests that refusals to harmful queries are mediated by a single “refusal direction” in activation space (Arditi et al., 2024), and that jailbreaks rely on manipulating this direction (Yu et al., 2024), yet these assumptions require further examination. Understanding refusal mechanisms is crucial, as it has direct implications for both offensive capabilities—informing more sophisticated jailbreaks (Huang et al., 2024)—and more importantly developing robust defensive strategies like improved adversarial training (Yu et al., 2024) and inference–time monitoring.
In this work, we go beyond extracting concepts using common input prompt methods by introducing a novel gradient-based approach to representation engineering which we use to investigate the mechanisms underlying refusal behavior in LLMs. We extract refusal-mediating directions more effectively, improving both precision and control while minimizing unintended side effects, which we demonstrate in Section 4. Unlike prior work that assumes model refusal is controlled by a single linear direction, we show in Section 5 that there exist multi-dimensional polyhedral cones which contain infinite refusal directions; we show an illustrative example in Figure 1. To further characterize refusal mechanisms in language models, we introduce representational independence, a criterion for identifying directions that remain mutually unaffected under intervention, capturing both linear and non-linear dependencies across layers. In Section 6, we demonstrate that even under this strict notion of independence, multiple complementary refusal directions exist. [^1] Resources & code: cs.cit.tum.de/daml/geometry-of-refusal
@@ 2 Background
Notation. Let $f:{\mathbb{T}}^{N_{\textnormal{seq}}}\to\Delta^{N_{\textnormal{seq}}\times|{\mathbb{T}}|}$ denote a language model, where $\Delta^{|{\mathbb{T}}|}$ is the probability simplex over vocabulary $\mathbb{T}$}. Given a prompt ${p}=(t_{1},\dots,t_{N_{\textnormal{seq}}})\in{\mathbb{T}}^{N_{\textnormal{seq}}}$ consisting of tokens $t_{i}$, each token is first embedded: $x_{i}^{(0)}=\textsc{Embed}(t_{i})$. The model then processes the token sequence through $L$ layers, where at each layer $l=1,\dots,L$ and token position $i$ the following transformation is applied:
@@
Extracting refusal directions. Paired prompts of harmful and harmless requests allow the extraction of a directional feature from the model’s residual stream as shown by prior work (Panickssery et al., 2024; Bolukbasi et al., 2016; Burns et al., 2024). Recent studies obtain this direction by computing the difference-in-means (DIM) (Panickssery et al., 2024; Arditi et al., 2024; Stolfo et al., 2024) between model representations on datasets of harmful prompts $\mathcal{D}_{\text{harm}}$ and harmless prompts $\mathcal{D}_{\text{good}}$:
@@
Here, ${\bm{x}}_{i}^{(l)}({p})$ represents the residual stream activations at position $i$, layer $l$ for input prompt p{p}.
Adversarial steering attacks. The extracted harmfulness direction can be used to manipulate the model’s refusal behavior. With white-box access, an attacker can prompt the model with harmful queries and suppress activations in the harmfulness direction, thereby reducing the model’s probability of refusal. This can be done through directional ablation of $\bm{r}$}} (where $^{\hat{{{\bm{r}}}}}$ denotes the unit vector) (Zou et al., 2023a):
@@
which subtracts a scaled $\bm{r}$}} from the residual stream.We follow common practice to apply both operations across all token positions and ablation across all layers while doing subtraction only at a single layer (Arditi et al., 2024).
@@ 3 Related Work
Adversarial attacks for LLMs. Many studies have explored hand–crafted adversarial techniques, such as persona modulation (Shah et al., 2023), language modifications (Zhu et al., 2023), or prompt engineering using repetitions and persuasive phrasing (Rao et al., 2024). Other works take a more systematic approach, employing techniques like genetic algorithms and random search (Chen et al., 2024), discrete optimization over input tokens (Zou et al., 2023b), or gradient-based methods to identify high-impact perturbations (Geisler et al., 2024). While identifying these vulnerabilities enables adversarial fine-tuning (Xhonneux et al., 2024) or improved training through Reinforcement Learning with Human Feedback (RLHF), recent works suggest that robustness remains a challenge (Zou et al., 2023a; Schwinn et al., 2024; Geisler et al., 2024; Scholten et al., 2025).
@@
Interpretability of LLMs. A parallel line of research focuses on understanding the internal mechanisms of LLMs, as their natural language outputs provide a unique opportunity to link internal states to interpretable behaviors. Interpretability research has led to the identification of various “features”—concepts represented by distinct activation patterns (Cunningham et al., 2023)—as well as “circuits”, which are subnetworks that implement a specific function or behavior. Prominent examples are backup circuits (Nanda et al., 2024) and information mover circuits (Wang et al., 2022). Many interpretability insights rely on extracting features using paired inputs with opposing semantics (Burns et al., 2024) and then manipulating residual stream activations to elicit specific behaviors (Panickssery et al., 2024). Representation engineering, as proposed by Zou et al. (2023a), investigates the linear representation of concepts such as truthfulness, honesty, and fairness in LLMs. The effectiveness of these methods supports the hypothesis that many features are encoded linearly in LLMs (Marks & Tegmark, 2024). These insights allow researchers to pinpoint and manipulate concept representations or specific circuits, enabling targeted debugging of behaviors, mitigating biases, and advancing safer, more reliable AI systems.
@@
Understanding Refusal Mechanisms. Recent research has focused on understanding the mechanisms underlying refusal behaviors in LLMs. For example, removing safety-critical neurons has been shown to decrease robustness (Wei et al., 2024; Li et al., 2024b). Zheng et al. (2024) demonstrate that adding explicit safety prompts shifts the internal representation along a harmfulness direction. O’Brien et al. (2024) propose to use sparse autoencoders to identify latent features that mediate refusal. The most relevant work to ours is Arditi et al. (2024), which builds on Zou et al. (2023a) and examines the representation of refusal in LLMs. Their work suggests that a single direction in a model’s activation space determines whether the model accepts or refuses a request. We challenge this notion by showing that refusal is mediated through more nuanced mechanisms. Concurrently, Pan et al. (2025) identify multiple independent refusal directions, providing more evidence to our findings in Section 6. While their focus is on representation shifts during safety finetuning, our work introduces a novel, gradient-based, top-down discovery approach applicable to any model.
Over-refusal Recent work addresses LLM over-refusal, where benign queries are rejected by overly strict safety filters. Röttger et al. (2023) (XSTest) and An et al. (2024) (PHTest) provide benchmarks to quantify false refusal rates, expanded by Cui et al. (2025)’s OR-Bench. Shi et al. (2024) attribute this to lexical shortcuts, proposing self-contrastive decoding for mitigation. Similarly, Li & Liu (2025) highlight over-defense in prompt injection guards with NotInject and InjecGuard. Collectively, these studies underscore the challenge of balancing safety and helpfulness, offering resources to diagnose and mitigate overrefusal.
4 Gradient–based Refusal Directions
@@
Research Question: Can our gradient–based representation engineering identify refusal directions?
@@
To investigate the refusal mechanisms in language models, we propose a gradient-based algorithm that identifies directions controlling refusal in the model’s activation space. We refer to it as Refusal Direction Optimization (RDO). Unlike prior approaches that extract refusal directions using paired prompts of harmless and harmful instructions (Arditi et al., 2024), our method leverages gradients to find better directions instead of solely relying on model activations. Similar to (Park et al., 2023), we define two key properties for refusal directions:
@@
  • Monotonic Scaling: when using the direction for activation addition/subtraction $\smash{\check{{\bm{x}}}^{(l)}_{i}={\bm{x}}^{(l)}_{i}+\alpha\cdot{{\bm{r}}}}$, the model’s probability of refusing instructions should scale monotonically with $\alpha$.
@@
We can encode the desired refusal properties into loss functions, allowing us to find corresponding refusal vectors $\bm{r}$}} using gradient descent. For the monotonic scaling property, we train the model to refuse harmless instructions ${p}_{\text{safe}}$} when running the model $f$ with a modified forward pass $f_{\text{add}({{\bm{r}}},l)}$ in which we add $\bm{r}$}} to the activations at layer $l$. We minimize the cross-entropy between the model output and target refusal response $t_{\text{refusal}}$}. For the surgical ablation property, we similarly compute the cross–entropy between a harmful response target $t_{\text{answer}}$} and the output of a modified forward pass $f_{\text{ablate}({{\bm{r}}})}$ to make the model respond to harmful instructions. A key strength of our gradient–based approach is the ability to control any predefined objective and thus we can control the extent to which other concepts are affected during interventions. For this, we use a retain loss based on the Kullback–Leibler (KL) divergence to ensure that directional ablation of $\bm{r}$}} on harmless instructions does not change the model’s output over a target response $t_{\text{retain}}$}. Algorithm 1 shows the full training procedure for our refusal directions.
@@
Setup. We construct a dataset of harmless and harmful prompts from the alpaca (Taori et al., 2023) and salad-bench (Li et al., 2024a) datasets (see Section A.1). An important consideration for our algorithm is the choice of targets $t_{\text{answer}}$} and $t_{\text{refusal}}$}. Generally, language models differ in their refusal and response styles, which is why we use model–specific targets rather than generating them via uncensored LLMs as in Zou et al. (2024). Specifically, we use the DIM refusal direction to generate our targets, though any effective attack can work. For the harmful answers $t_{\text{answer}}$}, we ablate the DIM direction and generate 30 tokens. Similarly, we use activation addition on harmless instructions to produce refusal targets $t_{\text{refusal}}$}. For helpful answers on harmless instructions that should be retained $t_{\text{retain}}$}, we generate 29 tokens without intervention. The retain loss $\mathcal{L}_{\text{retain}}$ is applied over the last 30 tokens, such that the last token of the model’s chat template is included. We detail hyperparameters and implementation in Appendix A.
@@
Evaluation. We evaluate our method by training a refusal direction on various models from the Gemma 2 (Team et al., 2024), Qwen2.5 (Yang et al., 2024), and Llama-3 (Dubey et al., 2024) families and compare against the DIM direction for which we use the same setup as Arditi et al. (2024) but with our expanded dataset. For a fair comparison, we train the refusal direction at the same layer that the DIM direction is extracted from, and during activation addition/subtraction set the scaling coefficient $\alpha$ to the norm of the DIM direction. We evaluate the jailbreak Attack Success Rate (ASR) on JailbreakBench (Chao et al., 2024) using the StrongREJECT fine–tuned judge (Souly et al., 2024). For inducing refusal via activation addition, we test 128 harmless instructions sampled from alpaca using substring matching of common refusal phrases. Model completions for evaluation are generated using greedy decoding with a maximum generation length of 512 tokens.
@@
Does the direction mediate refusal? In Figure 2, we show that for jailbreaking, our approach is competitive when using directional ablation and, on average, outperforms DIM when subtracting the refusal direction. Notably, despite not being explicitly optimized for subtraction–based attacks, our direction naturally generalizes to this setting. Figure 10 shows that adding the refusal direction to harmless inputs induces refusal more effectively with RDO than with DIM, further indicating that our method manipulates refusal more effectively.
@@
Is the direction more precise? To measure the side effects when intervening with the directions we track benchmark performance. Arditi et al. (2024) show that directional ablation with the DIM direction tends to have little impact on benchmark performance, except for TruthfulQA (Lin et al., 2021). In Table 1, we show that RDO impacts TruthfulQA performance much less severely, reducing the error by $40\%$ on average. We show the results for more benchmarks in Section B.3.
@@
We then evaluate the trade–off between safety and over–refusal for RDO and DIM on the XSTest benchmark (Röttger et al., 2023). As detailed in Appendix G, RDO consistently achieves a higher refusal rate for harmful inputs while maintaining or reducing the benign over–refusal rate compared to DIM. This means that for any given level of benign over–refusal, our method refuses more harmful requests, thereby yielding a uniformly better trade–off.
Is our method versatile? Hyperparameter tuning of the retain loss weight $\lambda_{\text{ret}}$ in Algorithm 1 allows for balancing between attack success and side effects (Section C.2). We observe that for many models—especially those in the Qwen 2.5 family—the majority of estimated DIM directions have too high side–effects, rendering it an unsuccessful attack (Figure 16). Our method is more flexible than previous work as we can choose the target layer freely while limiting side effects through the retain loss.
Key Takeaways. Our RDO yields more effective refusal directions with fewer side effects, establishing that gradient–based representation engineering is an effective approach for extracting meaningful directions, while allowing for more modeling freedom such as incorporating side constraints.